Find the mean of probability distribution of the random variable X, which can take only the values 2,4,5 and 9, given that P(2) = 9/20, P(4) = 1/20 , P(5) = 1/5 , and P(9) = 3/10


1
Expert's answer
2021-06-24T03:56:11-0400

"E(x)=\sum_{\forall x} xP(x)"

"=2(\frac{9}{20})+4(\frac{1}{20})+5(\frac{1}{5})+9(\frac{3}{10})"

"=4.8"